OpenStudy (anonymous):

pls help write three ratios equal to 4/36 i will medel!

OpenStudy (secret66):

36/4=9 1/9, 2/18, 3/27

OpenStudy (lanhikari22):

Did you know that if you multiply by any number on the upper side and the lower side, the number is equal to 1? For example, any number over itself is 1 (except for 0)\[\frac{8}{8}=\frac{cookies}{children}\] If you have 8 cookies, and 8 children, every child gets a cookie! The point is: Any number divided by itself gives you one, since division "divides" things into groups. \[\frac{9}{3}=\frac{cookies}{children}=\frac{9}{3}\cdot \frac{4}{4}=\frac{36}{12}\] You see if i had 4 times the number of children and 4 times the number of cookies, nothing changes! each child in this example will take 3 cookies, even if i had 1000 times the number of children, as long as ALSO have 1000 times the number of cookies! If you want to represent any fraction differently multiply the upper and the lower by the same amount/number.

OpenStudy (anonymous):

To find out your answers, just multiply 4/36 by any number (such as 2/2, 3/3, 4/4) and you will get an equivalent fraction.

OpenStudy (triciaal):

in general reduce the fraction to its lowest term then all multiples of that fraction will be equivalent. to get the multiples review what @LanHikari22 and @3lectric has above